OpenStudy (anonymous):

Can I have some help with Algebra 1?

OpenStudy (anonymous):

A system of equations is shown below: 6x = -7y + 11 5x + 2y = 13 What is the solution to this system of equations?

OpenStudy (anonymous):

Okay, So First you wanna isolate the X variable in the first equation.

OpenStudy (anonymous):

How would I do that? Sorry, I haven't done this in a while.

OpenStudy (anonymous):

@MarrBear ?

OpenStudy (anonymous):

6x+7y=11 5x+2y=13 --------------- 30x+35y=55 30x+10y=65 ------ 25y=-10 y=-10/25

OpenStudy (anonymous):

Okay, thank you

OpenStudy (anonymous):

Is that it?